OpenStudy (anonymous):

Solve x-19> *with the line under it* 23 A. x<*with the line under it*42 B. x>*with the line under it* 42 I choose B. Am I correct?

OpenStudy (lgbasallote):

yes...yes you are

OpenStudy (anonymous):

Thank you so much :D

OpenStudy (lgbasallote):

you only change symbols when dividing by a negative number..but when adding and subtracting both sides no change

OpenStudy (anonymous):

oh ok :D

OpenStudy (anonymous):

I'll have to remember that :)

OpenStudy (lgbasallote):

the way i remember it is this way...for example i have \[-x \le 4\] i'll make x positiveby adding x to both sides so i have \[0 \le 4 + x\] now i let x be alone by subtracting 4 from both sides \[-4 \le x\] in this method...it's the position of the left side and right side that interchanges and the symbol remains the same